OpenStudy (anonymous):

1+3+5+7+9....+97 gauss approach

OpenStudy (shubhamsrg):

gauss approach hmm, you know he is my favorite scientist! ^_^ so, what is the sum of first term and last term ?

OpenStudy (anonymous):

u add by to 1+2 is 3 3+2 is 5 and so on

OpenStudy (raden):

N = 1 + 3 + 5 + 7 + ... + 95 + 97 N = 97 + 95 + 93 + 91 + ... + 3 + 1 ------------------------------------ (+) 2N = 98 + 98 + 98 + 98 + ... + 98 + 98 2N = 98(49) N = 98(49)/2 = ...